# Blind and Topological Interference Managements for Bistatic Integrated Sensing and Communication

### v4

@@ II-B Performance Metrics
II-B1 Communication Degree of Freedom
To evaluate communication performance, we define cDoF as the sum degree of freedom per time slot, averaged over $t_{0}$ time slots. Each independent message symbol contributes one unit of communication DoF. The cDoF is computed as the total number of message symbols successfully decoded by all communication receivers, normalized by $t_{0}$. This metric quantifies the efficiency of communication in terms of the number of recoverable symbols over a given transmission duration.
II-B2 Sensing Degree of Freedom
To facilitate the characterization of the fundamental tradeoff between communication and sensing, we adopt a dual metric to quantify both functionalities. Thus we consider the sensing degree of freedom based on the number of effective (i.e., independent) observations of the target during the estimation and detection process as in [5, 36]. More specifically, assuming the total number of transmitting antennas connected to the sensor is ${\mathsf{m}}^{\prime}$, an effective observation sequence of length $N\geq{\mathsf{m}}^{\prime}$ implies that the transmitter emits a set of orthogonal sensing signals ${\mathbf{x}}_{1}^{[0]},{\mathbf{x}}_{2}^{[0]},\ldots,{\mathbf{x}}_{N}^{[0]}$ (each with dimension ${\mathsf{m}}^{\prime}\times 1$), satisfying $\begin{bmatrix}{\mathbf{x}}_{1}^{[0]},{\mathbf{x}}_{2}^{[0]},\ldots,{\mathbf{x}}_{N}^{[0]}\end{bmatrix}\begin{bmatrix}{\mathbf{x}}_{1}^{[0]},{\mathbf{x}}_{2}^{[0]},\ldots,{\mathbf{x}}_{N}^{[0]}\end{bmatrix}^{H}=I.$ Accordingly, as in [5, 36] the total sensing degree of freedom (sDoF) is defined as the total number of the independent observations $N$. Assume that the sensor obtains these $N$ independent observations by using $T$ time slots. The average sDoF (or just simply called sDoF in the rest) is $N/T$.
@@ III-A ISAC Model with Heterogeneous Coherence Times
For the bistatic ISAC systems with heterogeneous coherence times and interference channel including ${\mathsf{K}}$ single-antenna transmitters, ${\mathsf{K}}$ single-antenna communication receivers, and one sensor, the lower convex hull of the following tradeoff points is achievable:
@@
$(\text{sDoF},\text{cDoF})=(1,0)$, and $(\text{sDoF},\text{cDoF})=(0,1)$, where the latter two points are achieved by the existing sensing-only scheme and communication-only schemes.888 Note that owing to the delay requirement of ISAC systems (especially for the sensing tasks), we consider zero forcing for communications within a limited number of coherence blocks, which excludes the vanilla interference alignment strategies that achieve precise interference alignment at the cost of an extremely large block length [37]. Moreover, even compared with the time-sharing between $(\text{sDoF},\text{cDoF})=(1,0)$ and $(\text{sDoF},\text{cDoF})=(0,{\mathsf{K}}/2)$, when $\text{sDoF}=({\mathsf{K}}-1)/{\mathsf{K}}$ our scheme can achieve $\text{cDoF}=1$, while the above time-sharing can only achieve $\text{cDoF}=1/2$.
@@
For the bistatic ISAC systems with heterogeneous coherence times and the MU-MISO channel including one transmitter with ${\mathsf{m}}$ antennas, ${\mathsf{K}}$ single-antenna communication receivers, and one sensor, the lower convex hull of the following tradeoff points is achievable:
@@
For the bistatic ISAC systems with heterogeneous coherence times and the MU-MIMO channel including one transmitter with ${\mathsf{m}}$ antennas, ${\mathsf{K}}$ multi-antenna communication receivers (with the number of antennas ${\mathsf{n}}_{i},i=1,\ldots,{\mathsf{K}}$), and one sensor, the lower convex hull of the following tradeoff points is achievable:
@@ III-B ISAC Model with Heterogeneous Connectivity
For the bistatic ISAC model with heterogeneous connectivity described in Section II-A, Theorems 4 and Theorem 5 consider the non-cooperative and cooperative transmitters respectively, both for the Topology 1 and Topology 2.
Theorem 4 (Non-cooperative transmitters).
For the bistatic ISAC systems with heterogeneous connectivity and non-cooperative transmitters,
  • a.
    for the Topology 1 based on the $({\mathsf{K}}+1,U,D)$ neighboring antidotes network with ${\mathsf{K}}+1>U+D$ and $D\geq U$, the lower convex hull of the following trade-off points is achievable:
    $(\text{sDoF},\text{cDoF})=\left(\frac{1}{{\mathsf{K}}-D+U+1},\frac{{\mathsf{K}}(U+1)}{{\mathsf{K}}-D+U+1}\right),$ (4)
    $(\text{sDoF},\text{cDoF})=(1,0)$, and $(\text{sDoF},\text{cDoF})=\left(0,\frac{{\mathsf{K}}(U+1)}{{\mathsf{K}}-D+U+1}\right)$, where the latter two points are achieved by the existing sensing-only scheme and the communication-only scheme [25, 32];
  • b.
    for the Topology 2 based on the $({\mathsf{K}},U,D)$ neighboring antidotes network with ${\mathsf{K}}>U+D$ and $D\geq U$, if there exists one receiver $j\in\{1,\ldots,{\mathsf{K}}\}$ satisfying $\mathcal{R}_{s}=\mathcal{R}_{c}^{[j]}\setminus\mathcal{A}$ where $j\in\mathcal{A}$, the lower convex hull of the following trade-off points is achievable:
    $(\text{sDoF},\text{cDoF})=\left(\frac{1}{{\mathsf{K}}-D+U},\frac{{\mathsf{K}}(U+1)}{{\mathsf{K}}-D+U}\right),$ (5)
    $(\text{sDoF},\text{cDoF})=(1,0)$, and $(\text{sDoF},\text{cDoF})=\left(0,\frac{{\mathsf{K}}(U+1)}{{\mathsf{K}}-D+U}\right)$, where the latter two points are achieved by the existing sensing-only and communication-only schemes [25, 32].
Theorem 5 (Cooperative transmitters).
For the bistatic ISAC systems with heterogeneous connectivity and cooperative transmitters,
  • a.
    for the Topology 1 based on the $({\mathsf{K}}+1,d)$ regular network, the lower convex hull of the following trade-off points are achievable:
@@
    $(\text{sDoF},\text{cDoF})=(1,0)$, and $(\text{sDoF},\text{cDoF})=\left(0,\frac{2{\mathsf{K}}}{d+1}\right)$, where the latter two points are achieved by the existing sensing-only and communication-only schemes [26];
  • b.
    for the Topology 2 based on the $({\mathsf{K}},d)$ regular network, if there exists one receiver $j\in\{1,\ldots,{\mathsf{K}}\}$ satisfying $\mathcal{R}_{s}=\mathcal{R}_{c}^{[j]}\setminus\mathcal{A}$ where $j\in\mathcal{A}$, the lower convex hull of the following trade-off points is achievable:
For the bistatic ISAC systems with heterogeneous connectivity, compared to communication-only tradeoff points, in (4)-(7) we can maintain the same cDoFs while additionally obtaining non-zero sDoFs.
@@ IV Achievable schemes for the ISAC Model with Heterogeneous Coherence Times
This section provides the proposed ISAC schemes based on the BIA strategy for the bistatic ISAC system with heterogeneous coherence times, considering the interference and MU-MISO channels. The general description on the proposed scheme for the MU-MIMO channel (which extends the scheme for the MU-MISO channel) could be found in the Appendix A .
@@ IV-A Proof of Theorem 1: Interference Channel
Thus by removing the dedicated sensing signals (which are deterministic), the $k$-th communication receiver obtains from time slots $\{1,2,3\}$ :
@@
The sensing DoF is $2/3$, by obtaining $4$ effective observations can be obtained over $6$ time slots. Next, we demonstrate the signal design using blind interference alignment. Let ${\mathbf{x}}_{1}^{[0]}=[x_{1,1},x_{1,2},x_{1,3}]^{T}\in{\mathbb{C}}^{3\times 1}$, ${\mathbf{x}}_{2}^{[0]}=[x_{2,1},x_{2,2},x_{2,3}]^{T}\in{\mathbb{C}}^{3\times 1}$, ${\mathbf{x}}_{3}^{[0]}=[x_{3,1},x_{3,2},x_{3,3}]^{T}\in{\mathbb{C}}^{3\times 1}$, ${\mathbf{x}}_{4}^{[0]}=[x_{4,1},x_{4,2},x_{4,3}]^{T}\in{\mathbb{C}}^{3\times 1}$, satisfying $[{\mathbf{x}}_{1}^{[0]},{\mathbf{x}}_{2}^{[0]},{\mathbf{x}}_{3}^{[0]},{\mathbf{x}}_{4}^{[0]}][{\mathbf{x}}_{1}^{[0]},{\mathbf{x}}_{2}^{[0]},{\mathbf{x}}_{3}^{[0]},{\mathbf{x}}_{4}^{[0]}]^{H}=I$. Our idea is to use the fact that the channel for the sensor remains constant over a certain period and that the communication signals also remain constant to cancel the communication signals, which are interference for the sensor. For the $k$-th transmitter where $k\in\{1,2,3\}$, the transmitted signals over time slots $\{1,2,3\}$ are: $x^{[k]}(1)=W^{[k]}+x_{1,k}+x_{2,k}$, $x^{[k]}(2)=W^{[k]}+x_{2,k}$, $x^{[k]}(3)=W^{[k]}+x_{1,k}$. The signals received by the sensor can be expressed as:
@@
${\mathbf{y}}^{[k]}(t)=H^{[k1]}(t)({\mathbf{x}}_{c}^{[1]}(t)+{\mathbf{x}}_{s}^{[1]}(t))+H^{[k2]}(t)({\mathbf{x}}_{c}^{[2]}(t)$
$+{\mathbf{x}}_{s}^{[2]}(t))+\dots+H^{[k{\mathsf{K}}]}(t)({\mathbf{x}}_{c}^{[{\mathsf{K}}]}(t)+{\mathbf{x}}_{s}^{[{\mathsf{K}}]}(t))+{\mathbf{z}}^{[k]}(t),$
for each $t\in\{1,\ldots,2{\mathsf{K}}\}$. Since the CSI and the transmitted sensing signal $x_{s}^{[k]}(t)$ are known to the receivers, we can cancel this part of the signal at the receiver, resulting in an estimate $\hat{y}^{[k]}$ that only concerns the communication signals:
$\hat{y}^{[k]}(t)=H^{[k1]}(t)x_{c}^{[1]}(t)+\dots+H^{[k{\mathsf{K}}]}(t)x^{[{\mathsf{K}}]}(t)+z^{[k]}(t)$
$=H^{[k1]}(t)W_{1}^{[1]}+\dots+H^{[k{\mathsf{K}}]}(t)W_{1}^{[{\mathsf{K}}]}+z^{[k]}(t),$
for each $t\in\{1,\ldots,{\mathsf{K}}\}$. Since the channel coefficients $H^{[ki]}(t)$ where $t,i\in\{1,\ldots,{\mathsf{K}}\}$ are i.i.d. with circularly symmetric complex Gaussian distribution, the $k$-th receiver can decode $W_{1}^{[k]}$ from the first ${\mathsf{K}}$ time slots with high probability by using ZF. Similarly, it can decode $W_{2}^{[k]}$ from the second ${\mathsf{K}}$ time slots. Thus the cDoF is $1$.
Over $2{\mathsf{K}}$ time slots, the sensor can obtain $22{\mathsf{K}}-$ 2 effective observations, through $22{\mathsf{K}}-$ 2 sensing signals ${\mathbf{x}}_{k}^{[0]}=[x_{k,1},\ldots,x_{k,{\mathsf{K}}}]^{T}\in\mathbb{C}^{{\mathsf{K}}\times 1}$, where $k=1,2,\dots,2{\mathsf{K}}-2$, satisfying $[{\mathbf{x}}_{1}^{[0]},{\mathbf{x}}_{2}^{[0]},\dots,{\mathbf{x}}_{2{\mathsf{K}}-2}^{[0]}][{\mathbf{x}}_{1}^{[0]},{\mathbf{x}}_{2}^{[0]},\dots,{\mathbf{x}}_{2{\mathsf{K}}-2}^{[0]}]^{H}=I$.
@@
Recall that the communication message transmitted by each transmitter during the first ${\mathsf{K}}$ time slots remains the same, as well as the channel from each transmitter to the sensor. Thus for each $t\in\{2,3,\ldots,{\mathsf{K}}\}$, we have
$y_{s}^{[{\mathsf{K}}+1]}(1)-y_{s}^{[{\mathsf{K}}+1]}(t)=H^{[{\mathsf{K}}+1,1]}(1)x_{t-1,1}+$
$H^{[{\mathsf{K}}+1,2]}(1)x_{t-1,2}+\cdots+H^{[{\mathsf{K}}+1,{\mathsf{K}}]}(1)x_{t-1,{\mathsf{K}}}+z_{s}(1)-z_{s}(t)$
$=\begin{bmatrix}H^{[{\mathsf{K}}+1,1]}(1),\dots,H^{[{\mathsf{K}}+1,{\mathsf{K}}]}(1)\end{bmatrix}x_{t-1}^{[0]}+z_{s}(1)-z_{s}(t).$
Thus over the first ${\mathsf{K}}$ time slots, we obtain ${\mathsf{K}}-1$ effective observations through ${\mathbf{x}}_{1}^{[0]},\ldots,{\mathbf{x}}_{{\mathsf{K}}-1}^{[0]}$. Similarly, over the second ${\mathsf{K}}$ time slots, we can obtain another ${\mathsf{K}}-1$ effective observations through ${\mathbf{x}}_{{\mathsf{K}}}^{[0]},\ldots,{\mathbf{x}}_{2{\mathsf{K}}-2}^{[0]}$. Hence, the sensing degree of freedom (SDoF) is $({\mathsf{K}}-1)/{\mathsf{K}}$.

### v5

@@ II-B Performance Metrics
II-B1 Communication Degree of Freedom (cDoF)
To evaluate communication performance, we define cDoF as the sum degree of freedom per time slot, averaged over $t_{0}$ time slots. Each independent message symbol contributes one unit of communication DoF. The cDoF is computed as the total number of message symbols successfully decoded by all receivers, normalized by $t_{0}$. This metric quantifies the efficiency of communication in terms of the number of recoverable symbols over a given transmission duration.
II-B2 Sensing Degree of Freedom (sDoF)
To facilitate the characterization of the fundamental tradeoff between communication and sensing, we adopt a dual metric to quantify both functionalities. Thus we consider the sensing degree of freedom based on the number of effective (i.e., independent) observations of the target during the estimation and detection process as in [5, 39]. More specifically, assuming the total number of transmitting antennas connected to the sensor is ${\mathsf{m}}^{\prime}$, an effective observation sequence of length $N\geq{\mathsf{m}}^{\prime}$ implies that the transmitters emits a set of orthogonal deterministic training sensing signals ${\mathbf{x}}_{1}^{[0]},{\mathbf{x}}_{2}^{[0]},\ldots,{\mathbf{x}}_{N}^{[0]}$ (each with dimension ${\mathsf{m}}^{\prime}\times 1$), satisfying
$\begin{bmatrix}{\mathbf{x}}_{1}^{[0]},{\mathbf{x}}_{2}^{[0]},\ldots,{\mathbf{x}}_{N}^{[0]}\end{bmatrix}\begin{bmatrix}{\mathbf{x}}_{1}^{[0]},{\mathbf{x}}_{2}^{[0]},\ldots,{\mathbf{x}}_{N}^{[0]}\end{bmatrix}^{H}=\textbf{I}_{{\mathsf{m}}^{\prime}}.$ (1)
Accordingly, as in [5, 39] the total sensing degree of freedom (sDoF) is defined as the total number of the independent observations $N$. Assume that the sensor obtains these $N$ independent observations by using $T$ time slots. The average sDoF (or just simply called sDoF in the rest) is $N/T$888 Note that when the statistic CSI is known at the transmitters, the transmitted deterministic sensing signal can be $\begin{bmatrix}{\mathbf{x}}_{1}^{[0]},{\mathbf{x}}_{2}^{[0]},\ldots,{\mathbf{x}}_{N}^{[0]}\end{bmatrix}={\mathbf{W}}{\mathbf{S}}_{D}$ with deterministic training signal ${\mathbf{S}}_{D}\in\mathbb{C}^{{\mathsf{m}}^{\prime}\times N}$ satisfying ${\mathbf{S}}_{D}{\mathbf{S}}_{D}^{H}=\textbf{I}_{{\mathsf{m}}^{\prime}}$ and a precoding matrix $\mathbf{W}\in\mathbb{C}^{{\mathsf{m}}^{\prime}\times{\mathsf{m}}^{\prime}}$ optimized to minimize detection/estimation errors by using the statistic CSI. Thus (1) becomes $\begin{bmatrix}{\mathbf{x}}_{1}^{[0]},{\mathbf{x}}_{2}^{[0]},\ldots,{\mathbf{x}}_{N}^{[0]}\end{bmatrix}\begin{bmatrix}{\mathbf{x}}_{1}^{[0]},{\mathbf{x}}_{2}^{[0]},\ldots,{\mathbf{x}}_{N}^{[0]}\end{bmatrix}^{H}={\mathbf{W}}{\mathbf{W}}^{H}$. Except the modification on (1), the definition of sensing DoF remains the same. The proposed ISAC schemes in this paper can be directly extended according to this definition of sDoF, while the achieved tradeoff points between cDoF and sDoF do not change.
@@ III-A ISAC Model with Heterogeneous Coherence Times
For the bistatic ISAC systems with heterogeneous coherence times and interference channel including ${\mathsf{K}}$ single-antenna transmitters, ${\mathsf{K}}$ single-antenna receivers, and one sensor, the lower convex hull of the following tradeoff points is achievable:
@@
$(\text{sDoF},\text{cDoF})=(1,0)$, and $(\text{sDoF},\text{cDoF})=(0,1)$, where the latter two points are achieved by the existing sensing-only scheme and communication-only schemes. [^9] Note that owing to the delay requirement of ISAC systems (especially for the sensing tasks), we consider zero forcing for communications within a limited number of coherence blocks, which excludes the vanilla interference alignment strategies that achieve precise interference alignment at the cost of an extremely large block length [40]. Moreover, even compared with the time-sharing between $(\text{sDoF},\text{cDoF})=(1,0)$ and $(\text{sDoF},\text{cDoF})=(0,{\mathsf{K}}/2)$, when $\text{sDoF}=({\mathsf{K}}-1)/{\mathsf{K}}$ our scheme can achieve $\text{cDoF}=1$, while the above time-sharing can only achieve $\text{cDoF}=1/2$.
@@
For the bistatic ISAC systems with heterogeneous coherence times and the MU-MISO channel including one transmitter with ${\mathsf{m}}$ antennas, ${\mathsf{K}}$ single-antenna receivers, and one sensor, the lower convex hull of the following tradeoff points is achievable:
@@
For the bistatic ISAC systems with heterogeneous coherence times and the MU-MIMO channel including one transmitter with ${\mathsf{m}}$ antennas, ${\mathsf{K}}$ multi-antenna receivers (with the number of antennas ${\mathsf{n}}_{i},i=1,\ldots,{\mathsf{K}}$), and one sensor, the lower convex hull of the following tradeoff points is achievable:
@@
The description of the proposed ISAC schemes based on BIA for the interference channel and the MU-MISO channel can be found in Sections IV-A and IV-B respectively, while the ISAC scheme based on BIA for the MU-MIMO channel can be found in Appendix A.
@@ III-B ISAC Model with Heterogeneous Connectivity
For the bistatic ISAC model with heterogeneous connectivity described in Section II-A, Theorems 4 and 5 consider the two topology network $({\mathsf{K}}+1,U,D)$ neighboring antidotes network and $({\mathsf{K}}+1,d)$ regular network , respectively.
Theorem 4 ($({\mathsf{K}}+1,U,D)$ neighboring antidotes network).
For the bistatic ISAC systems with heterogeneous connectivity based on the $({\mathsf{K}}+1,U,D)$ neighboring antidotes network with ${\mathsf{K}}+1>U+D$ and $D\geq U$, the lower convex hull of the following trade-off points is achievable:
$(\text{sDoF},\text{cDoF})=\left(\frac{U+1}{{\mathsf{K}}-D+U+1},\frac{{\mathsf{K}}(U+1)}{{\mathsf{K}}-D+U+1}\right),$ (5)
$(\text{sDoF},\text{cDoF})=(1,0)$, and $(\text{sDoF},\text{cDoF})=\left(0,\frac{{\mathsf{K}}(U+1)}{{\mathsf{K}}-D+U+1}\right)$, where the latter two points are achieved by the existing sensing-only scheme and the communication-only scheme [28, 35]; ,
Theorem 5 ($({\mathsf{K}}+1,d)$ regular network).
For the bistatic ISAC systems with heterogeneous connectivity and cooperative transmitters, based on the $({\mathsf{K}}+1,d)$ regular network, the lower convex hull of the following trade-off points are achievable:
$(\text{sDoF},\text{cDoF})=\left(\frac{2}{d+1},\frac{2{\mathsf{K}}}{d+1}\right),$ (6)
The description of the proposed ISAC schemes based on TIM for the $({\mathsf{K}}+1,U,D)$ neighboring antidotes network and the $({\mathsf{K}}+1,d)$ regular network can be found in Sections V-A and V-B, respectively.
Remark 1 (Extended framework for the proposed schemes: Adding a sensor).
Given an existing topological network, another possible way to consider the sensor into the network is to purely add the sensor into the network, instead of replacing one receiver by this sensor as we considered previously. More precisely, in the $({\mathsf{K}},d)$ regular network or the $({\mathsf{K}},U,D)$ neighboring antidotes network, with the original ${\mathsf{K}}$ users being receivers and one additional user as the sensor, where the set of connected transmitters of the sensor is a subset of one receiver.
The proposed ISAC schemes for the $({\mathsf{K}}+1,d)$ regular network or the $({\mathsf{K}}+1,U,D)$ neighboring antidotes network could be directly extended to this framework. For the $({\mathsf{K}},U,D)$ neighboring antidotes network with an adding sensor, the lower convex hull of the following trade-off points is achievable:
$(\text{sDoF},\text{cDoF})=\left(\frac{U+1}{{\mathsf{K}}-D+U},\frac{{\mathsf{K}}(U+1)}{{\mathsf{K}}-D+U}\right),$ (7)
$(\text{sDoF},\text{cDoF})=(1,0)$, and $(\text{sDoF},\text{cDoF})=\left(0,\frac{{\mathsf{K}}(U+1)}{{\mathsf{K}}-D+U}\right)$, where the latter two points are achieved by the existing sensing-only scheme and communication-only schemes [28, 35]; the detailed description of the scheme can be found in the Appendix B-A.
For the $({\mathsf{K}},d)$ regular network with an adding sensor, the lower convex hull of the following trade-off points is achievable:
@@
$(\text{sDoF},\text{cDoF})=(1,0)$, and $(\text{sDoF},\text{cDoF})=\left(0,\frac{2{\mathsf{K}}}{d+1}\right)$, where the latter two points are achieved by the existing sensing-only and communication-only schemes [29]; the detailed description of the scheme can be found in the Appendix B-B.
@@ IV Achievable schemes for the ISAC Model with Heterogeneous Coherence Times
This section introduces the proposed BIA-based ISAC schemes for bistatic ISAC systems under heterogeneous coherence times, covering interference channel, MU-MISO, and MU-MIMO channels. The approach evolves from single-antenna to multi-antenna configurations, harnessing system diversity to effectively incorporate sensing capabilities at the sensor side.
@@ IV-A Proof of Theorem 1: Interference Channel
Thus by removing the dedicated sensing signals (which are deterministic), the $k$-th receiver obtains from time slots $\{1,2,3\}$ :
@@
The sensing DoF is $2/3$, by obtaining $4$ effective observations can be obtained over $6$ time slots. Next, we demonstrate the signal design using blind interference alignment. Let ${\mathbf{x}}_{k}^{[0]}=[x_{k,1},x_{k,2},x_{k,3}]^{T}\in\mathbb{C}^{3\times 1}$ for $k=1,2,3,4$, satisfying $[{\mathbf{x}}_{1}^{[0]},{\mathbf{x}}_{2}^{[0]},{\mathbf{x}}_{3}^{[0]},{\mathbf{x}}_{4}^{[0]}][{\mathbf{x}}_{1}^{[0]},{\mathbf{x}}_{2}^{[0]},{\mathbf{x}}_{3}^{[0]},{\mathbf{x}}_{4}^{[0]}]^{H}=\mathbf{I}_{3}$. Our idea is to use the fact that the channel for the sensor remains constant over a certain period and that the communication signals also remain constant to cancel the communication signals, which are interference for the sensor. For the $k$-th transmitter where $k\in\{1,2,3\}$, the transmitted signals over time slots $\{1,2,3\}$ are: $x^{[k]}(1)=W^{[k]}+x_{1,k}+x_{2,k}$, $x^{[k]}(2)=W^{[k]}+x_{2,k}$, $x^{[k]}(3)=W^{[k]}+x_{1,k}$. The signals received by the sensor can be expressed as:
@@
$y^{[k]}(t)=H^{[k1]}(t)(x_{c}^{[1]}(t)+x_{s}^{[1]}(t))+H^{[k2]}(t)(x_{c}^{[2]}(t)$
$+x_{s}^{[2]}(t))+\dots+H^{[k{\mathsf{K}}]}(t)(x_{c}^{[{\mathsf{K}}]}(t)+x_{s}^{[{\mathsf{K}}]}(t))+z^{[k]}(t),$
for each $t\in\{1,\ldots,2{\mathsf{K}}\}$. Since the CSI and the transmitted sensing signal $x_{s}^{[k]}(t)$ are known to the receivers, we can cancel this part of the signal at the receiver, resulting in an estimate $\hat{{\mathbf{y}}}^{[k]}$ that only concerns the communication signals in ${\mathsf{K}}$ time slots:
$\hat{{\mathbf{y}}}_{1}^{[k]}=\begin{bmatrix}\hat{y}^{[k]}(1)\\ \hat{y}^{[k]}(2)\\ \vdots\\ \hat{y}^{[k]}({\mathsf{K}})\end{bmatrix}=\sum_{i=1}^{{\mathsf{K}}}\begin{bmatrix}H^{[ki]}(1)x^{[k]}_{c}(1)\\ H^{[ki]}(2)x^{[k]}_{c}(2)\\ \vdots\\ H^{[ki]}({\mathsf{K}})x^{[k]}_{c}({\mathsf{K}})\end{bmatrix}+\begin{bmatrix}z^{[k]}(1)\\ z^{[k]}(2)\\ \vdots\\ z^{[k]}({\mathsf{K}})\end{bmatrix}$
$=\sum_{i=1}^{{\mathsf{K}}}\begin{bmatrix}H^{[ki]}(1)\\ H^{[ki]}(2)\\ \vdots\\ H^{[ki]}({\mathsf{K}})\end{bmatrix}W_{1}^{[i]}+\begin{bmatrix}z^{[k]}(1)\\ z^{[k]}(2)\\ \vdots\\ z^{[k]}({\mathsf{K}})\end{bmatrix}.$ (10)
Since the channel coefficients $H^{[ki]}(t)$ where $t,i\in\{1,\ldots,{\mathsf{K}}\}$ are i.i.d. with circularly symmetric complex Gaussian distribution, the $k$-th receiver can decode $W_{1}^{[k]}$ from the first ${\mathsf{K}}$ time slots with high probability by using ZF. Similarly, it can decode $W_{2}^{[k]}$ from the second ${\mathsf{K}}$ time slots. Thus the cDoF is $1$.
Over $2{\mathsf{K}}$ time slots, the sensor obtains $22{\mathsf{K}}-$ 2 effective observations by transmitting $22{\mathsf{K}}-$ 2 sensing signals ${\mathbf{x}}_{k}^{[0]}=[x_{k,1},\ldots,x_{k,{\mathsf{K}}}]^{T}\in\mathbb{C}^{{\mathsf{K}}\times 1}$, where $k=1,2,\ldots,2{\mathsf{K}}-2$, satisfying $[{\mathbf{x}}_{1}^{[0]},{\mathbf{x}}_{2}^{[0]},\dots,{\mathbf{x}}_{2{\mathsf{K}}-2}^{[0]}][{\mathbf{x}}_{1}^{[0]},{\mathbf{x}}_{2}^{[0]},\dots,{\mathbf{x}}_{2{\mathsf{K}}-2}^{[0]}]^{H}=\textbf{I}_{\mathsf{K}}$.
@@
Recall that the communication message transmitted by each transmitter during the first ${\mathsf{K}}$ time slots remains the same, as well as the channel from each transmitter to the sensor. Thus for each $\tau\in\{2,3,\ldots,{\mathsf{K}}\}$, we have
$\Delta y_{s}^{[{\mathsf{K}}+1]}(\tau)=y_{s}^{[{\mathsf{K}}+1]}(1)-y_{s}^{[{\mathsf{K}}+1]}(\tau)=H^{[{\mathsf{K}}+1,1]}(1)x_{\tau-1,1}+$
$H^{[{\mathsf{K}}+1,2]}(1)x_{\tau-1,2}+\cdots+H^{[{\mathsf{K}}+1,{\mathsf{K}}]}(1)x_{\tau-1,{\mathsf{K}}}+z_{s}(1)-z_{s}(\tau)$
$=\begin{bmatrix}H^{[{\mathsf{K}}+1,1]}(1),\dots,H^{[{\mathsf{K}}+1,{\mathsf{K}}]}(1)\end{bmatrix}{\mathbf{x}}_{\tau-1}^{[0]}+z_{s}(1)-z_{s}(\tau).$
Thus over the first ${\mathsf{K}}$ time slots, we obtain ${\mathsf{K}}-1$ effective observations through ${\mathbf{x}}_{1}^{[0]},\ldots,{\mathbf{x}}_{{\mathsf{K}}-1}^{[0]}$. Similarly, over the second ${\mathsf{K}}$ time slots, we can obtain another ${\mathsf{K}}-1$ effective observations through ${\mathbf{x}}_{{\mathsf{K}}}^{[0]},\ldots,{\mathbf{x}}_{2{\mathsf{K}}-2}^{[0]}$. Hence, the sDoF is $\frac{{\mathsf{K}}-1}{{\mathsf{K}}}$.